OpenStudy (anonymous):

How to factor 25a^1-(3a+2b)^2?

OpenStudy (mertsj):

25a^1 ????

OpenStudy (anonymous):

maybe \(25a^2\)?

OpenStudy (anonymous):

whoops yes i meant 25 squared!

OpenStudy (anonymous):

so the problem is 25a squared - (3a +2b) squared

OpenStudy (anonymous):

\[25a^2-(3a+2b)^2\] use \(x^2-y^2)=(x+y)(x-y)\) with \(x=5a,y=(3a+2b)\)

OpenStudy (anonymous):

That's the difference of two squares: so \[a^2-b^2=(a-b)(a+b) \] set \[a=\sqrt{25a^1} \] or \[a=\sqrt{25a^2} \] whichever one it is. Do the same for the second term, set b to equal the square root of the second term.

OpenStudy (anonymous):

YES. my book says factoring when either square is a square of a polynomial

OpenStudy (anonymous):

so satellite73, would the answer be (5a+3a+2b)(5a-3a+2b)??

OpenStudy (anonymous):

It works whether they are squares or not.

OpenStudy (anonymous):

i see, so would that be the answer?? or would you factor both of those trinomials and multiply?

OpenStudy (anonymous):

What you have is almost right: \[ [5a+(3a+2b)]*[5a-(3a+2b))\] \[ [5a+(3a+2b)]*[5a-3a-2b))\] check that -2b, you had "+2b" instead.

OpenStudy (anonymous):

i see, so when i multiply that, i should get my original question?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

alright, thank you!